OpenStudy (anonymous):

what is an equation for the line with slope 2/3 and y-intercept 9?

jigglypuff314 (jigglypuff314):

use slope-intercept form: y = mx + b when m = slope and b = y-intercept so plug in m = 2/3 and b = 9 into the form

OpenStudy (anonymous):

y=2/3x + 9

jigglypuff314 (jigglypuff314):

yep, so that's your equation :)